OpenStudy (anonymous):

Solve the equation for y. 4x – 5y = 9 A. y=4/5x-9/5 B. y=-4x-9 C. y=5/4x+9/4 D. y=-5/4x-5/9

OpenStudy (akashdeepdeb):

\[4x-5y = 9\] \[4x - 5y + 5y = 9 + 5y\] \[4x = 9 + 5y\] \[4x - 9 = 9 + 5y - 9\] \[4x-9 = 5y\] \[\frac{4x-9}{5} = \frac{5y}{5} = y\] \[y = \frac{4}{5} x - \frac{9}{5}\]